Shift Summary DAY Sept 14, 1999

Douglas A. Jensen (djensen@fnal.gov)
Tue, 14 Sep 1999 15:38:21 -0500 (CDT)

To: shiftsum@vivsgi.fnal.gov

Shift Summary for DAY shift, 09.14.99

On Shift: Emmanuel, DougJ

Hours of beam delivered at all intensities: 8.0 (/8.5)
Hours of beam delivered at nominal intensity: 8
Hours of E832 data (special runs not included): 7.5
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: assorted spills missing. Restart daq and
new run - some of beam rather spurious.

========================================================================
Runs ended this shift:

Tapes brought to FCC this shift:

========================================================================
Accesses during shift: none

========================================================================
Overview of shift: Mostly data taking E832 at 4E12. End with strange
DAQ hang - see below.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14518 30 07:04 inherit running exp. from owl shift
60 07:45 BOA = 1.23, not too bad. MCR called to say
positions may wander a bit. Tuning, Meson off
63 Beam off quite a bit. Data may be bad.
65 Spill back on target.
9:00 From time to time, get DA full messages. Checking
120 9:06 got a life-time=0 warning, then pipeline error
14518 127 9:16 end run. Do a fresh start. DAQ rather a mess.

14519 3 9:30 Running, after KFRIENDs etc. But beam lousy.
They are resoring beam to Meson.
15 9:45 Beam is now centered.
17 beam gone away. Linac power supply, no estimate.
30 10:02 Beam returns.
32 10:05 beam gone again
37 10:16 beam back again BOA 1.35 ( gets worse during spill,
e.g. 1.2 rising to 1.13 by the end - watching scaler)
53 10:39 beam off
58 10:43 beam on.
77 11:12 beam back after loosing 10 spills
157 13:00 DC Alcohol added by DN. Was 2.12, fill to about 3.2"
Question - will DC appearant gain reduce. Seems
effective gain is about 50 v too high ( in those
units )
13:20 Things settle down, Control room population drops
from 10+ down to the nominal 2.
200 14:40 Beam gone, MI controls problems. no est.
2?? 14:55 End run. Intend to update fera peds while the beam
is awol
> beam was off for ~8 spills, ending with 247.
247 15:00 Run did not stop. DAQ monitor screen froze, but
tapes running, triggers running
~255 END RUN

saved the histograms.
new tapes, fresh start, fera peds updated.

15520 15:17 Start the run
0 15:27 starting run with KFRIEND after fresh start.
3 15:30 shift ends back up to about `10 people here.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

Remember that reg HV is no longer interlocked to vacuum. If vacuum
fails, one must manually turn off reg HV. See note next to alarm screen
terminal.

Note that the air conditioning people will be here to start working
on things at 8 am. May loose air conditioning for a while.

Xerox still not working.